OpenStudy (anonymous):

please help! how many solutions does this have! either 1, none, or infinite. thanks in advance, will give best answer to person who explains best! a + 5 = 1/5 (15 + 5a) <--- Equation 1.) 1 2.) infinite 3.) none

OpenStudy (anonymous):

This would have 0 solutions. On the right hand side, you have \[\frac{ 1 }{ 5 }\left( 15+5a \right)\] which simplifies to \[3+a\] On the left hand side, you have \[a+5\]. Equating both sides of the equation you would have, \[3+a = a+5\] Bringing a over to the left hand side yields \[3+a-a=5\] Hence, the variable \[a\] gets cancelled out, and we say the equation has no solutions.

OpenStudy (anonymous):

thank you so much! thanks for showing work too, now I can solve them by myself!

OpenStudy (anonymous):

No problem at all. If you like an answer, please hit the Best Response button and close the question. :)